# RDF Stream Taxonomy: Systematizing RDF Stream Types in Research and Practice

### v2

@@ Abstract
Over the years, RDF streaming was explored in research and practice from many angles, resulting in a wide range of RDF stream definitions. This variety presents a major challenge in discussing and integrating streaming solutions, due to the lack of a common language. This work attempts to address this critical research gap, by systematizing RDF stream types present in the literature in a novel taxonomy. The proposed RDF Stream Taxonomy (RDF-STaX) is embodied in an OWL 2 DL ontology that follows the FAIR principles, making it readily applicable in practice. Extensive documentation and additional resources are provided, to foster the adoption of the ontology. Two realized use cases are presented, demonstrating the usefulness of the resource in discussing research works and annotating streaming datasets. Another result of this contribution is the novel nanopublications dataset, which serves as a collaborative, living state-of-the-art review of RDF streaming. The aim of RDF-STaX is to address a real need of the community for a better way to systematize and describe RDF streams. The resource is designed to help drive innovation in RDF streaming, by fostering scientific discussion, cooperation, and tool interoperability.
@@
RDF stream Taxonomy Ontology Nanopublications.
Resource type: Ontology, documentation, nanopublications dataset
DOI: 10.5281/zenodo.10072907
URL: https://w3id.org/stax/
@@ 1 Introduction
Over the years, the concept of RDF streams was investigated from many angles, resulting in streaming serializations, querying and reasoning engines, programming frameworks, datasets, and other resources [5, 50, 58, 60]. Without a doubt, streaming RDF solutions are in demand, and more innovation is still needed. However, the landscape of RDF streaming terminology is fragmented, with many dissimilar streaming task formulations and incompatible solutions.
@@
The term “RDF stream” is used in the literature to refer to many different concepts, ranging from theoretical models [11], to very technical ones, focusing on the representation and practical usage [17, 45, 50]. As noted in past research [14, 35, 49], RDF streaming systems need more standardization in stream exchange protocols and stream task definitions. At the same time, if the research is to solve real-world problems, the variety of the existing RDF streaming task formulations should be embraced, not limiting oneself to one specific type of RDF streams. However, up until now, there was very little academic discussion about the existing RDF stream types – what stream types are used, how they relate to each other, and which use cases are they applicable to. This paper aims to address this critical research gap.
@@
The main contribution of this work is the RDF Stream Taxonomy (RDF-STaX), embodied in an OWL 2 DL ontology. This novel taxonomy systematizes the existing RDF stream types on the basis of a broad state of the art review. RDF-STaX aims to be a practical and useful resource for researchers and practitioners, enabling them to semantically annotate RDF stream types in publications, software, and streams published on the Web. To achieve this, RDF-STaX is accompanied with extensive documentation, including usage examples. Finally, two implemented applications of RDF-STaX are presented – a living state-of-the-art review of RDF streaming using nanopublications, and annotating stream types in RiverBench datasets.
@@ 2 Literature Review
The primary goal of this paper is to introduce a taxonomy of RDF stream types, which naturally must be grounded in the state of the art to be relevant. As no comprehensive review of this topic is available, here we attempt to summarize the most significant works that use the concept of RDF streams.
The scope of this study only includes RDF streams that can be fully represented using the RDF 1.1 W3C Recommendation [10]. This notably excludes stream models that do not have a direct translation to RDF, such as RSP-QL [11]. However, due to the importance of such models, they are briefly discussed below. The reason behind this choice is that only streams fully representable in RDF can be easily exchanged and published on the Web, while the theoretical models have more limited applicability, e.g., only within one specific solution (e.g., a stream reasoning engine). As shown below, this scope choice still leaves a lot of variety to be explored.
The review was conducted by starting from several survey papers and books that focus on areas of research related to RDF streaming [5, 29, 39, 40, 43, 46, 53, 58, 60, 64], such as RDF data management, linked streaming data, and stream reasoning. Besides the works mentioned in these reviews, additional papers and software were added based on the authors’ expertise and by searching for keywords “RDF stream” and “RDF streaming” in Google Scholar. The review focused on works which use streams of RDF-based elements (triples, quads, graphs, datasets). For many of the examined works, determining the exact nature of the stream was not trivial, due to unclear descriptions and a lack of standardized terminology. Therefore, not only the papers were examined, but also the code and documentation, where available.
@@ 2.1 RDF Streaming Protocols
One area of research that mentions RDF streams are works focusing on streaming protocols for transmitting RDF data. One of the earlier examples is Streaming HDT (S-HDT) [26], which facilitates streaming RDF data in IoT devices. More specifically, producers (IoT devices) emit a stream of compressed, unnamed RDF graphs. RDSZ [18] is a later method that also attempts streaming compressed RDF graphs over the network. ERI [17] is a similar approach, however, it has several views of the RDF stream. At the highest level, ERI streams a continuous sequence of triples that is then divided into discrete blocks (RDF graphs). Then, on the lowest level, these graphs are decomposed into subject-molecules (graphs with a single RDF subject) that are then streamed over the network. Therefore, ERI has three different views of the streaming problem, with each corresponding to a different level in the protocol.
The RDF EXI communication protocol for constrained devices [28] focuses on streaming RDF graphs. Finally, the recent Jelly protocol [50] treats RDF streams similarly to ERI, with it streaming a sequence of RDF statements divided into discrete elements (stream frames). The main difference is that Jelly can also stream quad statements, making it generalizable to RDF datasets.
@@ 2.2 Stream Processing and Reasoning
There is a large body of research on the topics of RDF stream reasoning and stream processing, where a special focus is often placed on the time annotations of stream elements. We could find one work in this category that did not consider the temporal aspect, an early proposal for a streaming SPARQL engine from 2007 by Groppe et al. [23]. The proposed engine executes queries over streams of RDF triples.
@@
Most publications in this category focus on the inner workings of querying/reasoning engines, and adopt task formulations that are not fully grounded in the RDF specification. The most popular such formulation is the stream of timestamped triples [1, 3, 4, 8, 11, 13, 30, 34, 57, 64], where a stream is a sequence of $\langle triple,timestamp\rangle$ tuples. There are many variations of this task, with time intervals instead of timestamps [1], multiple timestamps per triple [12, 57], and different time domains. However, this model (and other equivalent or similar models, e.g., time-varying graphs [14, 33, 57]) by itself is not enough to fully realize an RDF stream – this would require encoding the time information in RDF, which is often omitted in these works.
@@
The lack of grounding in the RDF specification of the timestamped triple model is a known problem [8], and solutions to it were proposed. The most prominent formulation is the stream of timestamped named graphs, where each stream element is a named RDF graph. The timestamp (or other metadata) of this graph is added by making statements in the default graph about the named graph’s name node [7, 41, 63]. This representation was possibly first proposed by Tappolet & Bernstein in 2009 [56], and was then expanded by the W3C RSP Community Group, which created a draft specification for the RSP Data Model [47]. The RSP Data Model uses the aforementioned formulation to define RDF streams, but also mentions that RSP engines may consume streams of graphs, datasets, or named graphs and then turn them into RDF streams proper by associating each element with a timestamp. The RSP Data Model is the most mature and best-grounded model for representing timestamped streams, having been drafted by the community, and even proposing the formal semantics of RDF streams. However, it was never made into a W3C Recommendation, and the RSP Community Group now appears to be inactive [^1] https://lists.w3.org/Archives/Public/public-rsp/2023Mar/$0$.html.
@@
Time-Annotated RDF (TA-RDF) [48] is an alternative RDF representation for the timestamped stream problem (more precisely, TA-RDF has a direct translation to RDF). It identifies each stream element as an URI node in an RDF graph. The data relevant to that stream element is then attached to the node. This boils down to each stream element being an RDF graph in which a single node serves as the “point of entry”. This model is notably less flexible than the RSP Data Model. A similar approach is used by Stream Containers [49], a framework for publishing RDF streams on the Web. There, a stream is a sequence of unnamed RDF graphs that are referenced by their URIs. Each URI points to a node (the subject) within one of the graphs, and the temporal information is also contained within this graph.
@@
One auxiliary RDF stream definition also appears in this category. The VoCaLS vocabulary for describing linked streams [59] defines an RDF stream as a potentially infinite sequence of RDF graphs and/or triples. This definition is used in practice in the RDF4J API [57], whose YASPER implementation can consume streams of RDF graphs or RDF triples.
@@ 2.3 Streaming Semantic Annotation and Translation
Several systems were developed for streaming semantic annotation (lifting) and translation. RMLStreamer [24, 45] annotates data using the RML language [15], outputting a stream of RDF datasets. CARML [9] is a Java library that provides generic facilities for annotating data using RML. It supports outputting a stream of RDF quad statements (using RDF4J Rio [16]), which corresponds to a single RDF dataset. CARML does not natively support streams of datasets, and thus it operates on a different abstraction level than RMLStreamer. This limitation is addressed by the Semantic Annotation enabler [55] from ASSIST-IoT [54], which outputs a stream of RDF datasets. SPARQL-Generate [38] is an extension of the SPARQL language, which allows for generating RDF from various data sources. Its reference implementation [^2] https://github.com/sparql-generate/sparql-generate supports outputting a stream of RDF triples (using Apache Jena RIOT [2]).
Last in this category is the Inter Platform Semantic Mediator (IPSM) [20], which translates a stream of RDF graphs into another such stream, in real time. Although on a logical level IPSM processes RDF graphs, it outputs a stream of datasets, where each dataset has two named graphs – one containing the translated RDF, and the other with metadata about the translation itself.
@@ 2.4 Semantic Streaming Applications
Over the years, RDF streams were employed in various real-life use cases. One such example is DBpedia-Live [44] – a now-defunct service that published changes to the DBpedia knowledge base in real time. Two streams were published – deleted and added triples. In each stream, the elements were RDF graphs, in the form of N-Triples files.
RDF streams are also commonly used in semantic sensor networks [39]. In this context, the Graph of Things (GoT) [36] was proposed, which uses streams of RDF graphs. Another common use case is monitoring social media activity. Here, a survey by Keskisärkkä & Blomqvist [29] noted that RDF streams should have RDF graphs as elements. However, they also acknowledge that RDF streams can have RDF triples as elements. The recent proposal for a live open scientific knowledge graph [37] uses streams of RDF graphs to monitor scientific and social activity in real time.
@@ 2.5 Streaming I/O
The last broad category of RDF streaming solutions are those that process simple sequences of RDF statements (triples or quads) for input/output (I/O) operations. Two examples are the RDF4J Rio toolkit [16] and Apache Jena’s RIOT [2], both of which can parse and serialize streams of statements, reading from or writing to a Java byte stream. The two most obvious serializations for this task are N-Triples and N-Quads, however, Jena also supports streamed block formats for Turtle and Trig, where statements are transmitted in small batches (blocks).

### v3

@@ Abstract
Over the years, RDF streaming was explored in research and practice from many angles, resulting in a wide range of RDF stream definitions. This variety presents a major challenge in discussing and integrating streaming systems, due to the lack of a common language. This work attempts to address this critical research gap, by systematizing RDF stream types present in the literature in a novel taxonomy. The proposed RDF Stream Taxonomy (RDF-STaX) is embodied in an OWL 2 DL ontology that follows the FAIR principles, making it readily applicable in practice. Extensive documentation and additional resources are provided, to foster the adoption of the ontology. Three use cases for the ontology are presented with accompanying competency questions, demonstrating the usefulness of the resource. Additionally, this work introduces a novel nanopublications dataset, which serves as a collaborative, living state-of-the-art review of RDF streaming. The results of a multifaceted evaluation of the resource are presented, testing its logical validity, use case coverage, and adherence to the community’s best practices, while also comparing it to other works. RDF-STaX is expected to help drive innovation in RDF streaming, by fostering scientific discussion, cooperation, and tool interoperability.
@@
RDF stream; Taxonomy; Ontology; Interoperability; SKOS; Nanopublications.
[…]
@@ 1 Introduction
The Resource Description Framework (RDF) is a standard-backed foundation for modeling and exchanging knowledge graphs, with a wide range of practical applications Pan (2009); Hitzler (2021). The RDF model is based on the building blocks of the Web (e.g., it uses Internationalized Resource Identifiers for node names), making it very flexible and extensible Cyganiak et al. (2014). The list of use cases for RDF has expanded greatly since its inception, with time moving beyond applications where the modeled knowledge is treated as a single, static dataset. More recently, the concept of RDF streams has gained traction in research and industrial applications. Generally, an RDF stream refers to a potentially infinite stream of data modeled using the RDF standard, to which stream processing principles apply Kleppmann (2017). This research interest resulted in numerous contributions to the field, such as streaming serializations, querying and reasoning engines, programming frameworks, datasets, and other resources Bonte and Tommasini (2023); Sowiński et al. (2022); Tommasini et al. (2023); Van Assche et al. (2023). Without a doubt, streaming RDF solutions are in demand, and more innovation is still needed. However, the landscape of RDF streaming terminology is fragmented, with many dissimilar streaming task formulations and incompatible solutions.
@@
The term “RDF stream” is used in the literature to refer to many different concepts, ranging from theoretical models Dell’Aglio et al. (2014), to very technical ones, focusing on the representation and practical usage Fernández et al. (2014); Oo et al. (2022); Sowiński et al. (2022). As noted in past research works Dell’Aglio et al. (2016); Le-Phuoc et al. (2009); Schraudner and Harth (2022), RDF streaming systems need more standardization in stream exchange protocols and stream task definitions. This would be beneficial not only in the academic context, but would also facilitate the distribution and reuse of RDF streams in practice. Additionally, it could enable streaming tool interoperability through the inclusion of appropriate metadata. At the same time, if the research is to solve real-world problems, the variety of existing RDF streaming task formulations must be embraced, not limiting oneself to one specific type of RDF streams. However, up until now, there was very little academic discussion about the existing RDF stream types – what stream types are used, how they relate to each other, and which use cases are they applicable to. Our work aims to address this critical research gap, which has a large impact on the uptake of RDF streaming in general.
@@
The main contribution of this work is the RDF Stream Taxonomy (RDF-STaX), embodied in an OWL 2 DL ontology. This novel taxonomy systematizes the existing RDF stream types on the basis of a broad state-of-the-art review. RDF-STaX aims to be a practical and useful resource for researchers and practitioners, enabling them to semantically annotate RDF stream types in publications, software, and streams published on the Web. To achieve this, RDF-STaX follows the community’s best practices for ontology publishing and is accompanied with extensive documentation, including usage examples. The three envisioned use cases for RDF-STaX are presented with accompanying competency questions implemented in SPARQL. Two of these use cases were already implemented in practice – a living state-of-the-art review of RDF streaming using nanopublications, and annotating stream types in RiverBench datasets. Finally, RDF-STaX is compared to other ontologies, highlighting the gap which it covers.
This work is organized as follows. Section 2 presents a comprehensive literature review of the topic. Section 3 introduces the proposed systematization of RDF streams, along with an alignment of these definitions to the state of the art. Section 4 describes the proposed RDF-STaX ontology. Section 5 presents the use cases envisioned for RDF-STaX, while Section 6 includes a multifaceted evaluation of the proposed ontology, including a comparison to other ontologies. Finally, Section 7 discusses the results and outlines future work directions, with Section 8 presenting the concluding remarks.
@@ 2 Literature Review
The primary goal of this paper is to introduce a taxonomy of RDF stream types, which naturally must be grounded in the state of the art to be relevant. As no comprehensive and up-to-date review of this topic is available, here we attempt to summarize the most significant works that use the concept of RDF streams.
The scope of this study only includes RDF streams that can be fully represented using the RDF 1.1 W3C Recommendation Cyganiak et al. (2014). This notably excludes theoretical stream models that do not have a direct translation to RDF, such as RSP-QL Dell’Aglio et al. (2014). However, due to the importance of such models, they are briefly discussed to provide additional context. The reason to not consider purely theoretical models is that they are not directly comparable to models that have a fixed representation in RDF 1.1, as they operate on a different level of abstraction. Therefore, including both purely theoretical and fully representable RDF streams in the review would make the resulting taxonomy methodologically flawed. This scope decision is additionally motivated by the fact that only streams fully representable in RDF can be easily exchanged and published on the Web using already existing protocols, while the theoretical models have more limited applicability, for example, only within one specific system (e.g., a stream reasoning engine). This choice is thus natural due this work’s focus on solving the challenge of streaming tool interoperability and harmonizing the ways in which RDF streams are exchanged.
To cover the various areas of research related to RDF streaming, several survey papers and books were included in the review Bonte and Tommasini (2023); Keskisärkkä and Blomqvist (2013); Llanes et al. (2016); Ma et al. (2016); Modoni et al. (2014); Özsu (2016); Su et al. (2016); Tommasini et al. (2023); Van Assche et al. (2023); Zhang et al. (2021), on topics such as RDF data management, linked streaming data, and stream reasoning. For many of the examined works, determining the exact nature of the stream was not trivial, due to unclear descriptions and a lack of standardized terminology. Therefore, not only the papers were examined, but also the code and documentation, where available.
@@ 2.1 RDF Streaming Protocols
One prominent area of research that uses RDF streams are works focusing on streaming protocols for transmitting RDF data. One of the earlier examples is Streaming HDT (S-HDT) Hasemann et al. (2012), which facilitates streaming RDF data in IoT devices. More specifically, in S-HDT producers (IoT devices) emit streams of compressed, unnamed RDF graphs. RDSZ Fernández et al. (2014) is a later method that also attempts streaming compressed RDF graphs over the network. ERI Fernández et al. (2014) is a similar approach, however, it has several views of the RDF stream. At the highest level, ERI streams a continuous sequence of triples that is then divided into discrete blocks (RDF graphs). Then, on the lowest level, these graphs are decomposed into subject-molecules (graphs with a single RDF subject) that are then streamed over the network. Therefore, ERI has three different views of the streaming problem, with each corresponding to a different level in the protocol.
The RDF EXI communication protocol for constrained devices Käbisch et al. (2015) focuses on streaming RDF graphs in a very efficient manner. Finally, the recent Jelly protocol Sowiński et al. (2022) treats RDF streams similarly to ERI, with it streaming a sequence of RDF statements divided into discrete elements (stream frames). The main difference is that Jelly can also stream quad statements, making it generalizable to RDF datasets. It should be noted here that none of these works assume time annotations to be a necessary part of RDF streams, in contrast to the works described in the following subsection.
@@ 2.2 Stream Processing and Reasoning
There is a large body of research on the topics of RDF stream reasoning and stream processing, where a special focus is often placed on the time annotations of stream elements. We could find only one work in this category that did not consider the temporal aspect, an early proposal for a streaming SPARQL engine from 2007 by Groppe et al. Groppe et al. (2007). The proposed engine executes queries over streams of RDF triples, in real time.
@@
Most publications in this category focus on the inner workings of querying/reasoning engines, and adopt task formulations that are not fully grounded in the RDF specification. The most popular such formulation is the stream of timestamped triples Anicic et al. (2011); Barbieri et al. (2010); Bolles et al. (2008); Calbimonte and Corcho (2016); Dell’Aglio et al. (2014, 2015); Komazec et al. (2012); Le-Phuoc et al. (2011); Tommasini et al. (2021); Zhang et al. (2021), where a stream is a sequence of $\langle triple,timestamp\rangle$ tuples. There are many variations of this task, with time intervals instead of timestamps Anicic et al. (2011), multiple timestamps per triple Dell’Aglio et al. (2013); Tommasini et al. (2021), and different time domains. However, this model (and other equivalent or similar models, e.g., time-varying graphs Dell’Aglio et al. (2016); Le Phuoc et al. (2015); Tommasini et al. (2021)) by itself is not enough to fully realize and exchange an RDF stream – this would require encoding the time information in RDF, which is often omitted in these works.
@@
The lack of grounding in the RDF specification of the timestamped triple model is a known problem Calbimonte and Corcho (2016), and solutions to it were proposed. The most prominent formulation is the stream of timestamped named graphs, where each stream element is a named RDF graph. The timestamp (or other metadata) of this graph is added by making statements in the default graph about the named graph’s name node Calbimonte (2017); Mauri et al. (2016); Wu et al. (2022). This representation was possibly first proposed by Tappolet & Bernstein in 2009 Tappolet and Bernstein (2009), and was then expanded by the W3C RSP Community Group, which created a draft specification for the RSP Data Model RDF Stream Processing Community Group . The RSP Data Model uses the aforementioned formulation to define RDF streams, but also mentions that RSP engines may consume streams of graphs, datasets, or named graphs and then turn them into RDF streams proper by associating each element with a timestamp. The RSP Data Model is the most mature and best-grounded model for representing timestamped streams, having been drafted by the community, and even proposing the formal semantics of RDF streams. However, it was never made into a W3C Recommendation, and the RSP Community Group at the time of this work’s submission appears to be inactive (accessed on 4 June 2024).
A similar representation for timestamped streams was proposed in 2010 by Barbieri & Della Valle Barbieri and Valle (2010). In the proposal, stream elements are named RDF graphs (called instantaneous graphs). These graphs are linked together with additional metadata placed in an additional named graph dedicated to this purpose (called a stream graph). Each instantaneous graphs’ name node is associated with a timestamp triple using the sld:receivedAt property. This formulation differs from the RSP Data Model mostly by placing the metadata in a named graph instead of the default graph.
@@
Time-Annotated RDF (TA-RDF) Rodrıguez et al. (2009) is an alternative RDF representation for the timestamped stream problem (more precisely, TA-RDF has a direct translation to RDF). It identifies each stream element as an URI node in an RDF graph. The data relevant to that stream element is then attached to the node. This boils down to each stream element being an RDF graph in which a single node serves as the “point of entry”. This model is notably less flexible than the RSP Data Model, as it requires the entire content of each stream element to be connected to this single node. A similar approach is used by Stream Containers Schraudner and Harth (2022), a framework for publishing RDF streams on the Web. There, a stream is a sequence of unnamed RDF graphs that are referenced by URIs. This is not the same as RDF named graphs, as each URI instead points to a node (the element’s subject) within one of the unnamed graphs, and the temporal information is also contained within this graph.
@@
One auxiliary RDF stream definition also appears in this category. The VoCaLS vocabulary for describing linked streams Tommasini et al. (2018) defines an RDF stream as a potentially infinite sequence of RDF graphs and/or triples. This definition is used in practice in the RSP4J API Tommasini et al. (2021), whose YASPER implementation can consume streams of RDF graphs or RDF triples.
@@ 2.3 Streaming Semantic Annotation and Translation
Several systems were developed for streaming semantic annotation (lifting) and translation. RMLStreamer Haesendonck et al. (2019); Oo et al. (2022) annotates data using the RML language Dimou et al. (2014), outputting a stream of RDF datasets. CARML CARML contributors (2023) is a Java library that provides generic facilities for annotating data using RML. It supports outputting a stream of RDF quad statements (using RDF4J Rio Eclipse Foundation, Inc. (2023)), which corresponds to a single RDF dataset. CARML does not natively support streams of datasets, and thus it operates on a different abstraction level than RMLStreamer. This limitation is addressed by the Semantic Annotation enabler Szmeja (2023) from the ASSIST-IoT project Szmeja et al. (2023), which outputs a stream of RDF datasets, each corresponding to one input document. SPARQL-Generate Lefrançois et al. (2016) is an extension of the SPARQL language, which allows for generating RDF from various data sources. Its reference implementation (accessed on 4 June 2024) supports outputting a stream of RDF triples (using Apache Jena RIOT Apache Software Foundation (2023)).
Last in this category is the Inter Platform Semantic Mediator (IPSM) Ganzha et al. (2017), which translates a stream of RDF graphs into another such stream, in real time. Although on a logical level IPSM processes RDF graphs, it physically outputs a stream of RDF datasets, where each dataset has two named graphs – one containing the translated RDF, and the other with metadata about the translation itself.
@@ 2.4 Semantic Streaming Applications
Over the years, RDF streams were employed in various real-life use cases. One such example is DBpedia-Live Morsey et al. (2012) – a now-defunct service that published changes to the DBpedia knowledge base in real time. Two streams were published – deleted and added triples. In each stream, the elements were RDF graphs, in the form of N-Triples files.
RDF streams are also commonly used in semantic sensor networks Llanes et al. (2016). In this context, the Graph of Things (GoT) Le-Phuoc et al. (2016) was proposed, which uses streams of RDF graphs. Another common use case is monitoring social media activity. Here, a survey by Keskisärkkä & Blomqvist Keskisärkkä and Blomqvist (2013) noted that RDF streams should have RDF graphs as elements, as this approach better addresses the needs of their use case. However, they also acknowledge that RDF streams can sometimes have RDF triples as elements. Streams are also used for sending updates between decentralized social network servers, as in the ActivityPub protocol Tallon and Webber (2018) standardized by the W3C. In ActivityPub, an item of activity is described with an RDF graph (in JSON-LD). These items form ordered streams of RDF graphs which are then exchanged between the nodes. The Linked Data Event Streams (LDES) Van Lancker et al. (2021) proposal addresses publishing streams composed of elements, where each element is a fragment of an RDF graph, identified by a specific subject node. Finally, the recent proposal for a live open scientific knowledge graph Le-Tuan et al. (2022) uses streams of RDF graphs to monitor scientific and social activity in real time.
@@ 2.5 Streaming I/O
The last broad category of RDF streaming solutions are those that process simple sequences of RDF statements (triples or quads) for input/output (I/O) operations. Two examples are the RDF4J Rio toolkit Eclipse Foundation, Inc. (2023) and Apache Jena’s RIOT Apache Software Foundation (2023), both of which can parse and serialize streams of statements, reading from or writing to a Java byte stream. The two most obvious serializations for this task are N-Triples and N-Quads, however, Jena also supports streamed block formats for Turtle and Trig, where statements are transmitted in small batches (blocks).